# Choosing the Better Bandit Algorithm under Data Sharing: When Do A/B Experiments Work?

### v1

@@ 4 Simulation
[…]
@@ 4.1 Sign violation and preservation
(a) Running greedy with $\epsilon$-greedy
(b) Running greedy with UCB
@@ 4.2 Probability of correct comparison
Individual runJoint run with data sharing$\mathbb{P}(R(\mathcal{A}_{1})>R(\mathcal{A}_{2}))$$\mathbb{P}\left(R(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))>R(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))\right)$
(a) Comparing two $\epsilon$-greedy algorithms
Individual runJoint run with data sharing$\mathbb{P}(R(\mathcal{A}_{1})>R(\mathcal{A}_{2}))$$\mathbb{P}\left(R(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))>R(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))\right)$
(b) Comparing two UCB algorithms
@@ 4.3 Misspecified prior in Thompson sampling
We consider Thompson sampling with a prior size of $m=5$. Thompson sampling algorithm is described in Algorithm 3 in Appendix B. For a parameter $\gamma\in[0,1]$, we set the prior of arm $1$ (better arm) as $\textrm{Beta}(m(1-\gamma),m\gamma)$ and the prior of arm $2$ (worse arm) as $\textrm{Beta}(m\gamma,m(1-\gamma))$. Informally, a prior of $\textrm{Beta}(m\gamma,m(1-\gamma))$ can be thought of as having $m$ Bernoulli observations with $\gamma$ fraction of value $1$ and $(1-\gamma)$ fraction of value 00. Hence, a small value of $\gamma$ means that the algorithm has a stronger belief that arm $1$ is indeed the better arm. Each algorithm holds its own prior and the prior is not shared across the two algorithms.
@@
(a) Expected regret
Individual run Joint run with data sharing$\mathbb{P}(R(\mathcal{A}_{1})>R(\mathcal{A}_{2}))$$\mathbb{P}(R(\mathcal{A}_{1}\mid(\mathcal{A}_{1},\mathcal{A}_{2}))>R(\mathcal{A}_{2}\mid(\mathcal{A}_{1},\mathcal{A}_{2})))$
(b) Probability of comparison
@@
Figure 6 shows the regret of a single algorithm in an individual run, and the difference between the regrets of two algorithms in a joint run, in terms of both the expected regret and the probability of correct comparison. We again observe that the better algorithm when running individually remains better when running jointly in terms of both metrics; in other words, the comparison is preserved under data sharing. Interestingly, contrary to Figure 5LABEL:sub@float:stochastic_ucb from the previous section, we empirically observe that the comparison has a lower accuracy when the two algorithms run jointly than individually. We posit that the shared data helps the algorithm with the worse prior correct its misspecification more quickly, and hence decreases the performance gap between the two algorithms when running jointly.
@@ Acknowledgments
We thank Ramesh Johari, Yusha Liu, Ashwin Pananjady, and Chao Qin for helpful comments and discussions.
@@ Appendix A Additional results on the EXP3 algorithm
The EXP3 algorithm, presented in Algorithm 2 and denoted by EXP3, is originally designed for adversarial bandit problems [6], but also recently attracts attention in the stochastic bandit setting [46, 45]. By [45, Theorem 4], running the EXP3 algorithm individually yields the following regret bound.
@@
We consider a different setting of data sharing, where the greedy algorithm uses the data generated by the EXP3 algorithm, but the EXP3 algorithm does not use data from the greedy algorithm. This one-way data sharing setting is for technical convenience. Specifically, in the definition of the EXP3 algorithm, it requires not only the observed reward $Y_{t}$ but also the sampling probability $\widetilde{\rho}_{t}(k)$ associated with the selected arm $k$ in Line 2 of Algorithm 2. However, this sampling probability is not well-defined for the samples shared by the greedy algorithm. Hence, we assume for simplicity that the EXP3 algorithm only utilizes data generated by itself.
[…]
@@ C.1.1 Notation
[…]
@@
Throughout the proofs, we assume without loss of generality that arm $1$ is the best arm. Then $\Delta_{k}=\mu_{1}-\mu_{k}$ for each suboptimal arm $k\neq 1$. For the UCB algorithm, we denote by $c_{t,s}$ the length of the confident interval at timestep $t$ for an arm with $s$ samples as defined in (2). Since that $\lvert\mathcal{H}_{t}\rvert=2(t-1)$ because there are $2(t-1)$ pulls in total by both algorithms before timestep $t$, we have
@@ C.1.3 Properties under data sharing
Consider any $\alpha\in[0,1]$. When we jointly runs the UCB algorithm $\texttt{UCB}_{\alpha}$ with any other algorithm $\mathcal{A}$, the expected regret of $\texttt{UCB}_{\alpha}$ satisfies
@@
The proof of this proposition is provided in Section C.9.1. Intuitively, this result states that any additional samples do not hurt the regret rate of the UCB algorithm. We then establish a lower bound for the UCB algorithm under data sharing as follows.
@@
The proof of this proposition is provided in Section C.9.2. We also establish a lower bound for the $\epsilon$-greedy algorithm.
@@
[…]
@@ C.4 Proof of Theorem 2
The proof largely follows [21, Appendix C], with the difference that we consider an anytime UCB algorithm, whereas the UCB algorithm considered in [21] assumes the knowledge of the horizon $T$ and uses a factor of $\sqrt{\log T}$ instead of $\sqrt{\log t}$ for the confidence bound. In the anytime algorithm, the empirical mean of an arm no longer concentrates around the true mean when the timestep $t$ is small. However, our goal is weaker as we only need to show that the optimal arm is pulled $\Omega(T)$ times (as opposed to $T-o(T)$ times). Therefore, it suffices to show that the optimal arm is pulled $\Omega(T)$ times during the time interval $[T/2,T]$. We show that the empirical mean of each arm always concentrates around its true mean during this time interval $[T/2,T]$ with high probability. Conditional on this event, each suboptimal arm can only be pulled a limited number of times in this interval. As a result, the optimal arm must be pulled $\Omega(T)$ times with high probability.
@@
$\mathbb{P}\left(\mathcal{E}_{k}\right)$ $\leq\sum_{t=\lceil T/2\rceil}^{T}\mathbb{P}\left(\lvert\overline{X}_{k,N_{k,t}(\texttt{tot})}-\mu_{k}\rvert>c^{\alpha}_{t,N_{k,t}(\texttt{tot})}\right)$
@@ C.5 Proof of Theorem 3
We apply a similar argument as in [55, Thm 4.1]. We decompose term $M_{1}$ as:
@@
where in the summation $s$ goes up to $2(t-1)$, because there are two samples collected per timestep under data sharing. For any fixed $s\in\left\{0,\dots,2(T-1)\right\}$, we have the deterministic relation
@@ C.6 Proof of Theorem 4
We separately prove the upper and lower bounds for the $\epsilon$-greedy algorithm.
Lower bound for $\epsilon$-greedy.
[…]
Upper bound for $\epsilon$-greedy.
[…]
$\epsilon_{t}(\alpha_{1},C)=\min\left\{1,C/(2t-2)^{1-\alpha_{1}}\right\}.$
[…]
$\mathbb{P}\left(I_{t}(\texttt{$\epsilon$-grd}_{\alpha_{1},C})=k\right)$
$\qquad\leq\frac{\epsilon_{t}(\alpha_{1},C)}{K}+\mathbb{P}\left(\overline{X}_{k,N_{k,t}(\texttt{tot})}\geq\overline{X}_{1,N_{1,t}(\texttt{tot})}\right)$
$\qquad\leq\frac{C}{K(2t-2)^{1-\alpha_{1}}}+\underbrace{\mathbb{P}\left(\overline{X}_{k,N_{k,t}(\texttt{tot})}\geq\mu_{k}+\frac{\Delta_{k}}{2}\right)}_{M_{t}}+\underbrace{\mathbb{P}\left(\overline{X}_{1,N_{1,t}(\texttt{tot})}\leq\mu_{1}-\frac{\Delta_{k}}{2}\right)}_{M^{\prime}_{t}},$ (24)
[…]
Now we bound term $M_{t}$; a symmetric argument applies to term $M_{t}^{\prime}$. Consider any suboptimal arm $k\neq 1$. Recall that $N_{k,t}^{R}(\texttt{$\epsilon$-grd}_{\alpha_{1},C})$ denotes the number of times in which arm $k$ is chosen in exploration steps. We define its expectation by:
$\tau(t):=\mathbb{E}\left[N_{k,t}^{R}(\texttt{$\epsilon$-grd}_{\alpha_{1},C})\right]=\frac{1}{K}\sum_{i=K+1}^{t-1}\epsilon_{i}(\alpha_{1},C).$
The variance of $N_{k,t}^{R}(\texttt{$\epsilon$-grd}_{\alpha_{1},C})$ is:
By Bernstein’s inequality (8b) in 2, we have
We bound the term $M_{t}$ as:
We now provide upper and lower bounds for $\tau(t)$. Recall that $\epsilon_{t}=\min\left\{1,C/(2t-2)^{1-\alpha_{1}}\right\}\geq\min\left\{1,C/(2t)\right\}$. In the proof of Theorem 1, we have established the lower bound (11) that for any $t\geq(C+1)^{2}$,
$\tau(t)\geq\frac{C\log(t)}{4K}.$ (28a)
On the other hand, an upper bound for $\tau(t)$ is
$\tau(t)\leq\frac{1}{K}\sum_{i=K+1}^{t-1}\epsilon_{i}(\alpha_{1},C)\leq\frac{1}{K}\sum_{i=K+1}^{t-1}\frac{C}{(2i-2)^{1-\alpha_{1}}}\leq\frac{Ct^{\alpha_{1}}}{\alpha_{1}K}.$ (28b)
$M_{t}$ $\leq\frac{Ct^{\alpha_{1}}}{2\alpha K}t^{-\frac{C}{40K}}+\frac{2}{\Delta_{k}^{2}}e^{-\Delta_{k}^{2}(\frac{C\log t}{16K}-1)}$
$\leq\frac{C}{2\alpha K}\cdot\frac{1}{t^{3-\alpha_{1}}}+\frac{2e^{\Delta_{k}^{2}}}{\Delta_{k}^{2}}\cdot\frac{1}{t^{2}},$ (29)
where the last inequality is true by the assumption that $C\geq\max\left\{120K,\frac{32K}{\Delta_{\textrm{min}}^{2}}\right\}$, A similar argument yields the same upper bound for $M_{t}^{\prime}$.
Let us now first consider $0<\alpha_{1}\leq 1$. Plugging (29) back to (24), we have
$\mathbb{E}[N_{k}(\texttt{$\epsilon$-grd}_{\alpha_{1},C})]$
$\qquad=\sum_{t=1}^{T}\mathbb{P}\left(I_{t}(\texttt{$\epsilon$-grd}_{\alpha_{1},C})=k\right)$
$\qquad\leq(C+1)^{2}+\sum_{t=(C+1)^{2}}^{T}\mathbb{P}\left(I_{t}(\texttt{$\epsilon$-grd}_{\alpha_{1},C})=k\right)$
$\qquad\stackrel{{\scriptstyle\text{(i)}}}{{\leq}}(C+1)^{2}+\sum_{t=(C+1)^{2}}^{T}\frac{C}{K(2t-2)^{1-\alpha_{1}}}+\sum_{t=(C+1)^{2}}^{T}\left(M_{t}+M_{t}^{\prime}\right)$
$\qquad\stackrel{{\scriptstyle\text{(ii)}}}{{\leq}}(C+1)^{2}+\frac{C}{K}\sum_{t=(C+1)^{2}}^{T}\frac{1}{(2t-2)^{1-\alpha_{1}}}+\frac{C}{K\alpha_{1}}\sum_{t=(C+1)^{2}}^{T}\frac{1}{t^{3-\alpha_{1}}}+\frac{4e}{\Delta_{k}^{2}}\sum_{t=(C+1)^{2}}^{T}\frac{1}{t^{2}}$ (30)
$\qquad\leq(C+1)^{2}+\frac{CT^{\alpha_{1}}}{K\alpha_{1}}+\frac{C}{K\alpha_{1}}\cdot\frac{1}{\left(2-\alpha_{1}\right)}\frac{1}{\left((C+1)^{2}-1\right)^{2-\alpha_{1}}}+\frac{4e}{\Delta_{k}^{2}}\cdot\frac{1}{((C+1)^{2}-1)}$
$\qquad=O(T^{\alpha_{1}}),$
[…]
$\mathbb{E}\left[R_{T}(\texttt{$\epsilon$-grd}_{\alpha_{1},C}\mid(\texttt{$\epsilon$-grd}_{\alpha_{1},C},\texttt{$\epsilon$-grd}_{\alpha_{2},C}))\right]=O(T^{\alpha_{1}}),$
for $0<\alpha_{1}\leq 1$. Finally, when $\alpha_{1}=0$, we follow the same derivation with the difference that in (30), the second term becomes
$\frac{C}{K}\sum_{t=(C+1)^{2}}^{T}\frac{1}{(2t-2)}=O(\log T),$
C.7 Proof of Theorem 5
@@
$\mathbb{E}\left[R_{T}(\texttt{UCB}_{\alpha_{1}}\mid(\texttt{UCB}_{\alpha_{1}},\texttt{UCB}_{\alpha_{2}}))\right]=\begin{cases}O(\log T)&\text{if }\alpha=0,\\ O(T^{\alpha_{1}})&\text{if }0<\alpha\leq 1.\end{cases}$ (31)
completing the proof of the desired upper bound.

### v2

@@ 5 Simulation
We now present simulation studies that validate our theoretical results and explore additional settings, including using the probability of correct comparison as the performance metric and examining the impact of a misspecified prior. We consider $K=2$ arms for simplicity, where rewards are drawn from Bernoulli distributions with means $\mu_{1}=0.8$ and $\mu_{2}=0.2$. The horizon of the problem is $T=100$ unless otherwise specified. For the $\epsilon$-greedy algorithm $\texttt{$\epsilon$-grd}_{\alpha,C}$, we set its exploration constant to be $C=1$ for simplicity. In all simulations, each data point is based on 10,000 runs. Error bars (barely visible) indicate the standard error of the mean. All code for simulation is provided at https://github.com/jingyanw/bandit-data-sharing.
@@ 5.1 Sign violation and preservation
((a))
((b))
@@ 5.2 Probability of correct comparison
[…]
((a))
[…]
((b))
@@ 5.3 Misspecified prior in Thompson sampling
We consider Thompson sampling with a prior size of $m=5$. Thompson sampling algorithm is described in Algorithm 4 in Appendix B. For a parameter $\gamma\in[0,1]$, we set the prior of arm $1$ (better arm) as $\textrm{Beta}(m(1-\gamma),m\gamma)$ and the prior of arm $2$ (worse arm) as $\textrm{Beta}(m\gamma,m(1-\gamma))$. Informally, a prior of $\textrm{Beta}(m\gamma,m(1-\gamma))$ can be thought of as having $m$ Bernoulli observations with $\gamma$ fraction of value $1$ and $(1-\gamma)$ fraction of value 0. Hence, a small value of $\gamma$ means that the algorithm has a stronger belief that arm $1$ is indeed the better arm. Each algorithm holds its own prior and the prior is not shared across the two algorithms.
@@
((a))
[…]
((b))
@@
Figure 7 shows the regret of a single algorithm in an individual run, and the difference between the regrets of two algorithms in a joint run, in terms of both the expected regret and the probability of correct comparison. We again observe that the better algorithm when running individually remains better when running jointly in terms of both metrics; in other words, the comparison is preserved under data sharing. Interestingly, contrary to Figure 66(b) from the previous section, we empirically observe that the comparison has a lower accuracy when the two algorithms run jointly than individually. We posit that the shared data helps the algorithm with the worse prior correct its misspecification more quickly, and hence decreases the performance gap between the two algorithms when running jointly.
5.4 Detecting sign violation with ramp-up experiments
We now simulate ramp-up experiments for the regimes of sign preservation and sign violation. We vary the treatment probability $p\in[0,1]$ following a Bernoulli assignment in Algorithm 2. We take examples from the sign violation regime (running the greedy algorithm and the UCB algorithm jointly) and the sign preservation regime (running a pair of UCB algorithms), and examine whether our detection procedure from Section 4 behaves correctly.
((a))
((b))
Figure 8: Ramp-up experiments with data sharing. Varying the treatment probability $p$, we plot the average per-step regret of each algorithm under data sharing (top panel) and the total regret, summed over both algorithms and divided by $2T$ (bottom panel).
In Figure 8, we plot in the top panel the regret of each algorithm, normalized by the number of steps allocated to this algorithm, with control in blue and treatment in red. We also plot the total regret in the bottom panel, normalized by the total number of timesteps $2T$. We gradually increase the treatment probability $p$, where $p=0$ (left end) and $p=1$ (right end) correspond to running only the control or only the treatment algorithm, respectively. In both examples, the treatment algorithm appears to perform better in the experiment under data sharing. In Figure 8(a), we use the greedy algorithm and the UCB algorithm (with $\alpha=0$) to illustrate the regime of sign violation. We observe that while the greedy algorithm always incurs a lower regret throughout the ramp-up experiment, the total regret is not monotonic. Specifically, when the treatment probability becomes large (around $p=0.96$), the total regret increases drastically, triggering a warning according to the detection criterion (7). In contrast, Figure 8(b) illustrates a ramp-up experiment with $\texttt{UCB}_{0.5}$ (control) and $\texttt{UCB}_{0}$ (treatment) in the sign preservation regime. We observe a monotonically decreasing total regret where no warning is triggered: Increasing the probability of the treatment algorithm leads to a lower total regret. All these results are consistent with our theory from Section 4.
@@ Acknowledgments
We thank Ramesh Johari, Yusha Liu, Ashwin Pananjady, Aaditya Ramdas, and Chao Qin for helpful comments and discussions.
@@ Appendix A Additional results on the EXP3 algorithm
The EXP3 algorithm, presented in Algorithm 3 and denoted by EXP3, is originally designed for adversarial bandit problems [6], but also recently attracts attention in the stochastic bandit setting [53, 52]. By [52, Theorem 4], running the EXP3 algorithm individually yields the following regret bound.
@@
We consider a different setting of data sharing, where the greedy algorithm uses the data generated by the EXP3 algorithm, but the EXP3 algorithm does not use data from the greedy algorithm. This one-way data sharing setting is for technical convenience. Specifically, in the definition of the EXP3 algorithm, it requires not only the observed reward $Y_{t}$ but also the sampling probability $\widetilde{\rho}_{t}(k)$ associated with the selected arm $k$ in Line 3 of Algorithm 3. However, this sampling probability is not well-defined for the samples shared by the greedy algorithm. Hence, we assume for simplicity that the EXP3 algorithm only utilizes data generated by itself.
Under this one-way data sharing, we denote the regret of the greedy algorithm by $\mathbb{E}[R_{T}(\texttt{grd}\mid\texttt{EXP3}\rightarrow\texttt{grd})]$ to emphasize that only the EXP3 algorithm supplies data to the greedy algorithm. Since the greedy algorithm does not supply data to the EXP3 algorithm, we have $\mathbb{E}[R_{T}(\texttt{EXP3}\mid\texttt{EXP3}\rightarrow\texttt{grd})]=\mathbb{E}[R_{T}(\texttt{EXP3})]$. We now provide regret guarantees for the EXP3 algorithm and the greedy algorithm under one-way data sharing.
@@ C.1.1 Notation
We use the shorthand notation $N_{k}(\mathcal{A}\mid(\mathcal{A},\mathcal{A}^{\prime})):=N_{k,T+1}(\mathcal{A}\mid(\mathcal{A},\mathcal{A}^{\prime}))$ for the total number of pulls by algorithm $\mathcal{A}$, and likewise $N_{k}(\texttt{tot}\mid(\mathcal{A},\mathcal{A}^{\prime})):=N_{k,T+1}(\texttt{tot}\mid(\mathcal{A},\mathcal{A}^{\prime}))$ for the two algorithms in total. For the $\epsilon$-greedy algorithm, we call a timestep an “exploration step”, if the $\epsilon$-greedy algorithm samples an arm uniformly at random at this timestep (which happens with probability $\epsilon_{t}$ after the first $K$ timesteps). We denote by $N_{1,t}^{R}(\texttt{$\epsilon$-grd})$ the number of pulls by exploration steps of the $\epsilon$-greedy algorithm, or $N_{1,t}^{R}(\texttt{$\epsilon$-grd}\mid(\texttt{$\epsilon$-grd},\mathcal{A}))$ if the $\epsilon$-greedy algorithm runs jointly with another algorithm $\mathcal{A}$. We have the deterministic relation that for any arm $k$, timestep $t$, and algorithm $\mathcal{A}$,
$N_{k,t}^{R}(\texttt{$\epsilon$-grd})$ $\leq N_{k,t}(\texttt{$\epsilon$-grd})$
@@
Throughout the proofs, we assume without loss of generality that arm $1$ is the best arm. Then $\Delta_{k}=\mu_{1}-\mu_{k}$ for each suboptimal arm $k\neq 1$. For the UCB algorithm, we denote by $c_{t,s}$ the length of the confident interval at timestep $t$ for an arm with $s$ samples as defined in (2). Since that $\lvert\mathcal{H}_{t}\rvert=2(t-1)$ because there are 2(t1)2(t-1) pulls in total by both algorithms before timestep $t$, we have
@@ C.1.3 Properties under data sharing
Consider any $\alpha\in[0,1]$. When we jointly run the UCB algorithm $\texttt{UCB}_{\alpha}$ with any other algorithm $\mathcal{A}$, the expected regret of $\texttt{UCB}_{\alpha}$ satisfies
@@
The proof of this proposition is provided in Section C.9.1. Intuitively, this result states that any additional samples do not hurt the regret rate of the UCB algorithm.
We then establish a lower bound for the UCB algorithm under data sharing as follows.
@@
The proof of this proposition is provided in Section C.9.2.
We also establish upper and lower bounds for the $\epsilon$-greedy algorithm:
@@
Consider any $\alpha\in[0,1]$ and any $C\geq\max\{120K,\,32K/\Delta_{\textrm{min}}^{2}\}$. When we jointly run the $\epsilon$-greedy algorithm $\texttt{$\epsilon$-grd}_{\alpha,C}$ with any other algorithm $\mathcal{A}$, the expected regret of $\texttt{$\epsilon$-grd}_{\alpha}$ satisfies
$\mathbb{E}\left[R_{T}(\texttt{$\epsilon$-grd}_{\alpha}\mid(\mathcal{A},\texttt{$\epsilon$-grd}_{\alpha}))\right]=\begin{cases}O(\log T)&\text{if }\alpha=0,\\ O(T^{\alpha})&\text{if }0<\alpha\leq 1.\end{cases}$
The proof of this proposition is provided in Section C.9.3.
Proposition 6.
Consider any $\alpha\in[0,1]$ and any $C>0$. When we jointly run the $\epsilon$-greedy algorithm $\texttt{$\epsilon$-grd}_{\alpha,C}$ with any other algorithm $\mathcal{A}$, then the expected regret of $\texttt{$\epsilon$-grd}_{\alpha,C}$ satisfies
@@ C.4 Proof of Theorem 2
The proof largely follows [22, Appendix C], with the difference that we consider an anytime UCB algorithm, whereas the UCB algorithm considered in [22] assumes the knowledge of the horizon $T$ and uses a factor of $\sqrt{\log T}$ instead of $\sqrt{\log t}$ for the confidence bound. In the anytime algorithm, the empirical mean of an arm no longer concentrates around the true mean when the timestep $t$ is small. However, our goal is weaker as we only need to show that the optimal arm is pulled $\Omega(T)$ times (as opposed to To(T)T-o(T) times). Therefore, it suffices to show that the optimal arm is pulled $\Omega(T)$ times during the time interval $[T/2,T]$. We show that the empirical mean of each arm always concentrates around its true mean during this time interval $[T/2,T]$ with high probability. Conditional on this event, each suboptimal arm can only be pulled a limited number of times in this interval. As a result, the optimal arm must be pulled $\Omega(T)$ times with high probability.
@@
[…]
@@ C.5 Proof of Theorem 3
We apply a similar argument as in [67, Thm 4.1]. We decompose term $M_{1}$ as:
@@
where in the summation $s$ goes up to 2(t1)2(t-1), because there are two samples collected per timestep under data sharing. For any fixed $s\in\left\{0,\dots,2(T-1)\right\}$, we have the deterministic relation
@@ C.6 Proof of Theorem 4
Upper bound for UCB.
@@
$\mathbb{E}\left[R_{T}(\texttt{UCB}_{\alpha_{1}}\mid(\texttt{UCB}_{\alpha_{1}},\texttt{UCB}_{\alpha_{2}}))\right]=\begin{cases}O(\log T)&\text{if }\alpha=0,\\ O(T^{\alpha_{1}})&\text{if }0<\alpha\leq 1,\end{cases}$ (28)
proving the desired upper bound for $\texttt{UCB}_{\alpha_{1}}$. A symmetric argument yields the upper bound for $\texttt{UCB}_{\alpha_{2}}$.
Lower bound for UCB.
@@
proving the desired lower bound.
Upper bound for $\epsilon$-greedy.
Applying Proposition 5 to $\texttt{$\epsilon$-grd}_{\alpha_{1},C}$ with the other algorithm $\mathcal{A}$ being $\texttt{$\epsilon$-grd}_{\alpha_{2},C}$ yields the desired upper bound for $\texttt{$\epsilon$-grd}_{\alpha_{1},C}$. A symmetric argument yields the upper bound for $\texttt{$\epsilon$-grd}_{\alpha_{2},C}$.
Lower bound for $\epsilon$-greedy.
Applying Proposition 6 to $\texttt{$\epsilon$-grd}_{\alpha_{2},C}$ with the other algorithm $\mathcal{A}$ being $\texttt{$\epsilon$-grd}_{\alpha_{1},C}$ yields the desired lower bound for $\texttt{$\epsilon$-grd}_{\alpha_{2},C}$.
C.7 Proofs of the detection procedure
In this section, we present the proofs of the theoretical results from Section 4 for the detection procedure.
C.7.1 Proof of Theorem 5
Lower bound.
The $\Omega(T^{\max\{\alpha_{1},\alpha_{2}+\beta-1\}})$ lower bound for the cases without $\alpha_{1}=1$ or $1-\beta\geq\alpha_{2}$. Denote by $\epsilon_{1,t}$ and $\epsilon_{2,t}$ the exploration probabilities of the two algorithms. The total number of pulls on any suboptimal arm $k$ is:
$\mathbb{E}[N_{k,T}(\texttt{tot})]$ $\geq\frac{1}{K}\sum_{t=1}^{2T}\big[(1-p)\cdot\epsilon_{1,t}+p\cdot\epsilon_{2,t}\big]$
$=\frac{1}{K}\sum_{t=1}^{2T}\left((1-p)\cdot\min\left\{1,\frac{C}{t^{1-\alpha_{1}}}\right\}+p\cdot\min\left\{1,\frac{C}{t^{1-\alpha_{2}}}\right\}\right).$
With a similar analysis to the proof of Proposition 6 in Section C.9.4, we have that for $0<\alpha_{1}\leq 1$,
$\mathbb{E}[N_{k,T}(\texttt{tot})]$ $\geq(1-p)\cdot\Omega(T^{\alpha_{1}})+p\cdot\Omega(T^{\alpha_{2}}),$
where the term $\Omega(T^{\alpha_{1}})$ is replaced by $\Omega(\log T)$ when $\alpha_{1}=0$. Setting $p=T^{\beta-1}$ and $T\geq 2^{\frac{1}{1-\beta}}$, we have $1-p=1-T^{\beta-1}>\frac{1}{2}$. Therefore, when $0<\alpha_{1}\leq 1$ or $\alpha_{2}+\beta>1$,
$\mathbb{E}[N_{k,T}(\texttt{tot})]$ $=\Omega(T^{\alpha_{1}})+\Omega(T^{\alpha_{2}+\beta-1}),$
and otherwise
$\mathbb{E}[N_{k,T}(\texttt{tot})]$ $=\Omega(\log T),$